OpenStudy (anonymous):

The graph of the following system of equations is 2x + y = 6 6x + 3y = 12

OpenStudy (anonymous):

A Overlapping lines B Parallel lines C Intersecting lines

hero (hero):

Hint: Divide both sides of the second equation by 3

OpenStudy (anonymous):

If you want to know that you should write both lines in the y = mx + b form. y = -2x + 6 y = (12 - 6x ) / 3 = -2x + 4 As you see both line has the -2 in front of the x. Therefore they are parallel to each other .

OpenStudy (anonymous):

Since the b term differ, 4 and 6, the lines are not overlapping :)

hero (hero):

2x + y = 6 2x + y = 4 Is sufficient enough to see that they are parallel lines.

hero (hero):

Since everything on the left side is the same.

OpenStudy (anonymous):

Interesting, thanks.

OpenStudy (anonymous):

thank you all for your help :)